TheSmartOne (thesmartone):

If a car's mass is 600 kg, and its speed is 10 m/s. And a scooter's mass is 80 kg, and the scooter's speed is 50 m/s. Then which will require more force to stop it if the distance is the same?

TheSmartOne (thesmartone):

@Michele_Laino :)

TheSmartOne (thesmartone):

I thought we used f = ma And then f = 600 x 10 f = 6,000 And f = 80 x 50 f = 4,000 So it would take more force to stop the car. Am I correct? :P

OpenStudy (michele_laino):

the object which requires more force is the object which has the highest kinetic energy

OpenStudy (michele_laino):

now the kinetic energy of the car is: 0.5 * 600*100=30,000 Joules whereas kinetic energy of scooter is: 0.5*80*2500=100,000 Joules

OpenStudy (michele_laino):

so it is the scooter

OpenStudy (michele_laino):

oops.. formula for kinetic energy is: \(KE=1/2 \cdot m \cdot v^2\)

TheSmartOne (thesmartone):

oh thank you so much :D

TheSmartOne (thesmartone):

@eta @waterineyes there is your answer :p